All question parts

Determine the equivalent resistance of the
Answer #1

a)

equivalent resistance starting from right side

R,R and R are in series

R+R+R=3R

3R and R are in parallel

3R*R(3R+R) =0.75R

0.75R ,R and R are in series

0.75R+R+R=2.75R

2.75R and R are in parallel

2.75R*R/(2.75R+R) =0.7333R

0.7333R ,R and R are in series

Req =R+R+0.7333R

Req =2.7333R

Req =2.7333*215

Req =587.67 ohms

b)

I=V/Req =50/587.67

I=0.085 A

c)

Ib =I*(2R/2R+0.7333R)

Ib=0.062 A